OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 5x − 6y = 6 x − y = 12 Add the two equations together Multiply the second equation by −6 Multiply the first equation by −6 Multiply the second equation by 6

OpenStudy (anonymous):

multiply the second equation by -6

OpenStudy (anonymous):

We have just done 3 of these..duh.

OpenStudy (amistre64):

5 -6 = 6 1 -1 = 12 y = (60-6)/(-5+6) = 54 / 1 x = (-72+6)/(-6+5) = 66/ 1 66 − 54 = 12 ... its good i believe

OpenStudy (amistre64):

i think that is called cramers rule that i did